OpenStudy (anonymous):

Solve the function step-by-step. Then Does the sequence converge, or diverge? (-1)^(n-1)/sqrt(n)

OpenStudy (anonymous):

@halorazer Any ideas?...Im done..My brain is gone :/

OpenStudy (zzr0ck3r):

what does solve a function mean?

OpenStudy (zzr0ck3r):

\(\frac{(-1)^{n-1}}{\sqrt{n}}\)?

OpenStudy (anonymous):

It just means(show the steps) solve it. lol

OpenStudy (anonymous):

And yes that is how the problem is set up. :)

OpenStudy (zzr0ck3r):

\(-1\le(-1)^{n-1}\le 1\) so \(\frac{-1}{\sqrt{n}}\le\frac{(-1)^{n-1}}{\sqrt{n}}\le\frac{1}{\sqrt{n}}\) now let \(n\rightarrow \infty\) \(0\le\frac{(-1)^{n-1}}{\sqrt{n}}\le 0\) so \(\lim_{n\rightarrow \infty}\frac{(-1)^{n-1}}{\sqrt{n}}=0\)

OpenStudy (anonymous):

So basically it converges?

OpenStudy (zzr0ck3r):

if you are asking that question I have a feeling you are confused. which part are you confused about.

OpenStudy (zzr0ck3r):

yes it converges to 0

OpenStudy (anonymous):

Thank you so much! :) It makes sense now, I got an answer similar but I couldn't figure out why 0 converged...I was making such a big(small) mistake. Haha now I can rest. :P